# Self-lensing of moving gravitational-wave sources can break the microlensing crossing timescale degeneracy

### v1

@@ IV.1 Joint observation of $d_{\rm LS}$ and $M_{{\rm MBH},z}$
$M_{{\rm MBH},z}=10^{4}\,M_{\odot}\,\left(\frac{\rm Hz}{f}\right)~,$ (18)
the limitation comes from being able to obtain $d_{\rm LS}$ from the crossing timescale, Eq. (13), shown in a red dashed line in Fig. 5. For a lower $M_{{\rm MBH},z}$ than Eq. (18), the limitation comes from being able to see at least a few modulations, Eq. (17), in the blue region in Fig. 5. For this latter case, the exact limiting curve depends on the various parameters in Eq. (17), and in general it is more restrictive on the maximum value of $d_{\rm LS}$ than Eq. (13).
@@
Figure 6: Summary of the relations between parameters and how to obtain $d_{\rm LS}$ and $M_{{\rm MBH},z}$ from measuring $t_{\rm E}$, $f$ and $T$. By measuring $t_{\rm E}$ from the width of the curve, we obtain $d_{\rm LS}$ through Eq. (2). In GWs, the interference pattern can provide information about the modulation period $T$. By observing $T$, the GW frequency $f$ and $t_{\rm E}$, we can estimate $M_{{\rm MBH},z}$ with Eq. (16).
The expected MBH masses and distance distributions depend on the astrophysical environment (star clusters, galactic nuclei, AGN disks) and its formation channels for GW sources. Measuring $d_{\rm LS}$ and $M_{{\rm MBH},z}$ from the same signal may help distinguish the environment where the GW signal comes from, by comparing them to the expected values of $d_{\rm LS}$ and $M_{{\rm MBH},z}$ for each environment.
@@ V Inferring the orbital inclination of the binary
It may be possible to infer the inclination of the orbit $\theta$ with respect to the line of sight if the distance $d_{\rm LS}$, the MBH mass $M_{\rm MBH}$, and $y_{0}$ [^4] The value of $y_{0}$ can be estimated from the relative amplitude of the curve, see Fig. 4. can be obtained from the signal. Given that $y_{0}\simeq\eta_{0}/R_{\rm E}$ is the normalization of the physical offset $\eta_{0}$ in self-lensing, and the inclination angle of the orbit with respect to the line of sight is given by $\theta\approx\eta_{0}/d_{\rm LS}$ for small angles, the inclination can be estimated as
@@ VII Conclusions
  • If the distance $d_{\rm LS}$, the mass $M_{\rm MBH}$ and $y_{0}$ could be simultaneously obtained in GWs, it would be possible to infer the inclination of orbit with respect to the line of sight, $\theta\propto y_{0}M_{\rm MBH}^{1/2}d_{\rm LS}^{-1/2}$ [Eq. (20)].
Obtaining $d_{\rm LS}$ and $M_{{\rm MBH},z}$ simultaneously would provide valuable information about the environment of the binary [e.g., samsing-25-AGN]. Consequently, it would help constrain the astrophysical environments where GW signals come from.

### v2

@@ IV.1 Joint observation of $d_{\rm LS}$ and $M_{{\rm MBH},z}$
$M_{{\rm MBH},z}\sim 10^{4}\,M_{\odot}\,\left(\frac{\rm Hz}{f}\right)~,$ (18)
the limitation comes from being able to obtain $d_{\rm LS}$ from the crossing timescale, Eq. (13), shown in a red dashed line in Fig. 5. For a lower $M_{{\rm MBH},z}$ than Eq. (18), the limitation comes from being able to see at least a few modulations, Eq. (17), in the blue region in Fig. 5. For the latter case, the exact limiting curve depends on the various parameters in Eq. (17), and in general it is more restrictive on the maximum value of $d_{\rm LS}$ than Eq. (13).
@@
Figure 6: Summary of the relations between parameters and how to obtain $d_{\rm LS}$ and $M_{{\rm MBH},z}$ from measuring $t_{\rm E}$, $f$ and $T$. By measuring $t_{\rm E}$ from the width of the curve, we obtain $d_{\rm LS}$ through Eq. (2). In GWs, the interference pattern can provide information about the modulation period $T$. By observing $T$, $f$ and $t_{\rm E}$, we can estimate $M_{{\rm MBH},z}$ with Eq. (16).
The expected MBH masses and distance distributions depend on the specific astrophysical environment (star cluster, galactic nucleus, AGN disk) and its formation channels for GW sources. Measuring $d_{\rm LS}$ and $M_{{\rm MBH},z}$ from the same signal may help distinguish the environment where the GW signal comes from, by comparing them to the expected values of $d_{\rm LS}$ and $M_{{\rm MBH},z}$ for each environment.
@@ V Inferring the orbital inclination of the binary
It may be possible to infer the inclination of the orbit $\theta$ with respect to the line of sight if the distance $d_{\rm LS}$, the MBH mass $M_{\rm MBH}$, and $y_{0}$ [^4] The value of $y_{0}$ can be estimated from the relative amplitude of the curve, see Fig. 4. can be obtained from the signal. Given that $y_{0}\simeq\eta_{0}/R_{\rm E}$ is the normalized physical offset in self-lensing, and the inclination angle of the orbit with respect to the line of sight is given by $\theta\approx\eta_{0}/d_{\rm LS}$ for small angles, the inclination can be estimated as
@@ VII Conclusions
  • If the distance $d_{\rm LS}$, the mass $M_{\rm MBH}$ and $y_{0}$ could be simultaneously obtained in GWs, it would be possible to infer the inclination of the orbit with respect to the line of sight, $\theta\propto y_{0}M_{\rm MBH}^{1/2}d_{\rm LS}^{-1/2}$ [Eq. (20)].
Obtaining $d_{\rm LS}$ and $M_{{\rm MBH},z}$ simultaneously would provide valuable information about the environment around the binary merger [e.g., samsing-25-AGN]. Consequently, it would help constrain the astrophysical environments where GW signals come from.